por q esta ecuacion 2x+8=-2x-6
Answer by stanbon(75887)

2x+8=-2x-6
Add 2x to both sides to get:
4x + 8 = -6
---
Subtract 8 from both sides:
4x = -14
----
Divide by 4 to get:
x = -7/2
==================
Cheers,
Stan H.
=================